Circular No. 8914

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2007 VQ_{11} (CATALINA) A slightly diffuse object found by A. Boattini on Catalina images taken

on 2008 Feb. 1.1 UT was posted on the Minor Planet Center’s ‘NEOCP’ webpage. J. Young (Table Mountain 0.61-m Cassegrain reflector) confirmed the object on CCD images from Feb. 1.2 UT as having a very diffuse, round, 8′′ coma with no tail visible in fair seeing with light cirrus clouds. T. Spahr then identified this object with several observations from 2007 Nov. and Dec., including an apparently asteroidal object discovered by the Catalina Sky Survey on Nov. 3 (discovery observation tabulated below) that had been designated 2007 VQ_{11} after being linked to observations made on Nov. 6 at the Xu-Yi station of Purple Mountain Observatory (cf. MPS 225184).

2007 UT  α_{2000}  δ_{2000}  Mag.
Nov. 3.42085  4^{h}11^{m}40.^{s}45  +3^{o}21′51′′.5  18.8

The new astrometry, the following orbital elements, and an ephemeris appear on MPEC 2008-C02.

Epoch = 2008 Feb. 24.0 TT

T = 2008 Feb. 13.5568 TT  ω = 277.^{o}6529
e = 0.502035  Ω = 164.0574 }2000.0
q = 2.693808 AU  i = 12.3249

a = 5.409639 AU n^{o} = 0.0783342 P = 12.582 years

COMETS C/2007 U5 AND C/2007 U6 (SOHO) Further to IAUC 8912, additional Kreutz-sungrazing comets have been

found on SOHO website images, as tabulated below. C/2007 U5 appeared stellar (mag ∼ 7.5) in C3 images, and was slightly diffuse with no tail in C2 images. C/2007 U6 was very faint (mag ∼ 8) and quite diffuse.

Comet 2007 UT α_{2000} δ_{2000} Inst. F MPEC
C/2007 U5 Oct. 24.488 13^{h}45.^{m}8 −13^{o}04′ C3/2 MU 2008-A51
C/2007 U6  24.729 13 49.2 −12 55 C2  AK 2008-A51

COMET 17P/HOLMES Corrigendum. Regarding the gas-expansion velocity reported on IAUC

8891, the correct value is ∼ 1.0 km/s, further confirmed using the CS(3-2) line.

2008 February 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green